OpenStudy (anonymous):

For the remaining three problems, write always, sometimes, or never. If it is always true, give the property or two examples. If it is never true, give the property or two examples. If it is sometimes true, give one example for which it is true and one example for which it is false. 1. x + x = 2x 2. x • x = 2x 3. x^2 = -9

OpenStudy (perl):

1. always true 2. sometimes true . (only when x = 0, 2 ) 3. never true (unless you are using non-real numbers)

OpenStudy (carson889):

1. Always true ie. 3 + 3 = 2*6 --> 6 = 6 True, 4 + 4 = 2*4 -->8 = 8 True. 2. Sometimes, only two values make it true. ie: 0,2 3. Never true, because when you multiply any two of the same real numbers (negative or positive) together they equal a positive.